OpenStudy (anonymous):

Hello Openstudy! I was hoping to find some help on homework problems I am working on that I must turn in by the end of the night tonight, would anyone be so kind as to help me with these problems and explain them to me? if so I would greatly appreciate it and would fan and metal anyone willing to help. Thank you! -Krizzzy46

OpenStudy (anonymous):

Details: They are statistics and algebra 2 problems mostly

OpenStudy (triciaal):

nice letter! as soon as you post I will try to help with the algebra questions

OpenStudy (anonymous):

Thank you!, one moment please

OpenStudy (anonymous):

A company manufactures 2,000 units of its flagship product in a day. The quality control department takes a random sample of 40 units to test for quality. The product is put through a wear-and-tear test to determine the number of days it can last. If the product has a lifespan of less than 26 days, it is considered defective. The table gives the sample data that a quality control manager collected. 39 31 38 40 29 32 33 39 35 32 32 27 30 31 27 30 29 34 36 25 30 32 38 35 40 29 32 31 26 26 32 26 30 40 32 39 37 25 29 34 The point estimate of the mean of the sample is__________ , and the point proportion of the defective units is___________

OpenStudy (triciaal):

what is the range? (highest - lowest) to average divide by 2 arrange all the numbers from lowest to highest the middle is the median = good estimate the real mean is the sum of all divided by 40

OpenStudy (triciaal):

less than 26 days = defective from the group of 40 how many are less than 26? ratio defective = no <26 / 40

OpenStudy (triciaal):

did this have any more sub-parts? example how many are expected to be bad for the day? 1/20*2000 expected 100 defects

OpenStudy (anonymous):

2 are less than 26, does that mean I would divide 40 by 2? and no It didn't

OpenStudy (triciaal):

defect rate = 2/40 = 1/20

OpenStudy (anonymous):

Is there a specific formula I would follow to find the point estimate mean and the point proportion of the defective units?

OpenStudy (triciaal):

what is the point of telling what the sample size is ?

OpenStudy (triciaal):

then again my stats is not current maybe those new terms have more to it than what I learnt

OpenStudy (triciaal):

@satellite73 please help (real question posted )

OpenStudy (anonymous):

no the program they are having me use has some questionable... questions haha. They arent very descriptive